a bomber is flying horizontally with a constant speed of 150m/s at a height of 78.4m. the pilot has to drop a bomb at the enemy target. at what horizontal distance from the target should he release the bomb so that it hits the target

Grade:11

3 Answers

Divyansh Nema
5 years ago
  Let h be horizontal distance ,
  
  G = 9.8 m/s^2
  U = 150 m/s
  H = 78.4 m 
 
 
  h  =  u ( sq.root 2H / g  )
 
  h  = 150 ( 2 )    =  300 m 
  
Manda Snehashish Reddy
5 years ago
u = 150 m/s
H = 78.4 m/s
g = 9.8 m/s2
 
We know that R=u(2H/g)1/2
                     R=150 x (2 x 78.4/9.8)1/2
                     R=150 x (16)½
                     R=600 metres
Manda Snehashish Reddy
5 years ago
Nema, you forgot to consider another 2 somewhere, check your answer again. No disrespect here, just telling you